OpenStudy (anonymous):

what is the surface area of a cube if the volume is 125 cubic in.?

OpenStudy (anonymous):

x^3=125

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

area of each face = 5*5=25 cube has six face so 25*6=